Calculate point elasticity of demand for demand function Q=10-2p for
decrease in price from Rs 3 to Rs 2
1
Expert's answer
2014-04-04T13:44:06-0400
Ed = (deltaQ/Q0)/(deltaP/P0) = (deltaQ/deltaP)/(P0/Q0)
Where,
deltaQ = variation of the demand
deltaP = variation of the price
Q0 = quantity demanded at time t0
P0 = price at time t0
If Q = 10 - 2p, P0 = Rs 3, P1 = Rs 2, then Q0 = 10 - 2*3 = 4 units, Q1 = 10 - 2*2 = 6 units
Ed = ((6 - 4)/4)/((2 - 3)/3) = -0.5/(1/3) = -1.5, so the demand is elastic